6) Objects with charges +Q and +4Q are situated as shown above. At which point could...

Z

6) Objects with charges +Q and +4Q are situated as shown above. At which point could you place a third object of charge +Q so that it would be in equilibrium?

A) A

B) B

C) C

D) D

E) E
Answer #1

let us consider a section of the unit length

since both the charges are positive so null point must be in between these two charges

then,

at equilibrium point electric field due to +Q and +4Q is equal

k Q / d^2 = k (4 Q ) / (6-d)^2

( ( 6 - d) / d) ^2 = 4Q / Q

(6 - d) / d = sqrt (4)

6/d - d/d = sqrt (4)

6/d - 1 = sqrt (4)

d = 2

so D is the point where net electric field is supposed to be zero.

hence option (D) is correct